A hyperthermic (feverish) male, with a body mass of 104 kg, has a mean body temperature of 107F. He is to


be cooled to 98.6F by placing him in a water bath, which is initially at 77F. Calculate what is the minimum


volume of water required to achieve this result. The specific heat capacity of a human body is 3.5 kJ/(kg·K). The


specific heat capacity for water is 4186 J/(kg·K). You must first find an appropriate formula, before substituting


the applicable numbers.

1
Expert's answer
2022-05-25T10:27:10-0400

"c_1m_1(t_1-t)=c_2m_2(t-t_2),"

"m_2=m_1\cdot\frac{c_1}{c_2}\cdot\frac{t_1-t}{t-t_2}=33.8~kg,"

"V=33.8~l."